75percent of a first order reaction was completed in 32 minutes.when was 50percent of the reaction completed?
1
Expert's answer
2012-07-17T09:06:55-0400
v=
dC/dt.
v=const
v1=v2
v1=dC1/dt1
v2=dC2/dt2
dC1/dt1=dC2/dt2
(75)/32=(50)/x
2.34375=50/x
2.34375x=50
x=21.33
min